# FedMLAC: Mutual Learning Driven Heterogeneous Federated Audio Classification

### v1

@@ 4.1 FedMLAC Overview
The second component, FedMLAC-Aggregation, operates on the server side. It aggregates the updated Plug-in AC models uploaded by the clients, resulting in an updated global Plug-in AC model. This updated global AC model serves as an auxiliary tool for the subsequent round of client AC model updates, as illustrated in Fig. 1.
@@ 4.2 FedMLAC Update
3:     for each epoch $i$ from $1$ to $E$ do
@@
5:              Compute loss $\mathcal{L}^{k}(\mathcal{B};\omega^{k}_{t},\theta_{t}^{k})$ using Eq. (7)
@@
10:              Compute loss $\mathcal{L}^{g}(\mathcal{B};\omega^{k}_{t},\theta^{k}_{t})$ using Eq. (8)
@@
4.2.1 Local AC Model Update
@@
wherein $KL(\cdot,\cdot)$ denotes the KL divergence loss. By combining Eqs. (5) and (6), we reformulate the client AC model training objective function in the $(t+1)^{th}$ communication round as:
@@
4.2.2 Global Plug-in Model Update
@@
This loss ensures that the Plug-in model effectively integrates client-specific AC data knowledge, enhancing its adaptability and robustness for global FL. The entire FedMLAC-Update process is detailed in Algorithm 1.
4.3 FedMLAC Aggregation
Once the global Plug-in model on each client completes its local knowledge update, it is uploaded to the server for aggregation to generate a new global Plug-in model. In traditional federated aggregation (e.g., Eq. (3)), all client-uploaded Plug-in models $\{\theta_{t+1}^{k}\}_{k\in S_{t}}$ are directly combined using data-size weighted averaging. However, under non-IID data distributions or adversarial conditions (e.g., noisy inputs, label-flipping, or poisoned data), some local Plug-in models may introduce substantial deviations from the true global optimum. These deviations manifest as abnormally large or small parameters in specific layers, thereby biasing the global Plug-in model when such updates are indiscriminately aggregated.
@@ 4.3 FedMLAC Aggregation
This value reflects how much Plug-in model $k$’s layer-$l$ parameters deviate from the average behavior of the cohort. A high deviation may indicate either statistical heterogeneity or an adversarial anomaly.
@@
4:         for each model layer $l$ from $1$ to $L$ do
5:              Compute $\bar{\theta}_{t+1}^{l}$ using Eq. (9)
@@
7:              Compute $\delta^{k,l}$ using Eq. (10)
@@
12:     Sort and prune $\bar{\mathcal{T}}_{l}$ as $\mathcal{T}_{l}$ according to Eq. (11)
@@
To mitigate the influence of outlier updates, we prune the Plug-in models that exhibit extreme deviations at each layer. Specifically, for each layer $l$, we first compute the deviation scores $\delta^{k,l}$ for all clients $k\in S_{t}$ using Eq. (10). These scores are then sorted in ascending order, and both the top-$\lfloor v_{h}\cdot|S_{t}|\rfloor$ and bottom-$\lfloor v_{l}\cdot|S_{t}|\rfloor$ clients are removed from consideration. The remaining clients form the trusted subset $\mathcal{T}_{l}$ used for layer-$l$ aggregation, formally defined as:
@@
4.4 Disscusion
FedMLAC embodies a unified design philosophy that bridges personalization, scalability, and robustness in FedAC. Its core strength lies in decoupling global knowledge aggregation from personalized model learning through a collaborative Plug-in mechanism. This architectural separation not only supports model heterogeneity, but also allows clients to retain autonomy over their model structures and local objectives—crucial in diverse AC scenarios where tasks, devices, and environments differ significantly.
Moreover, by framing mutual learning as bidirectional knowledge distillation, FedMLAC integrates both generalization and adaptation within each training round. This mitigates the classic trade-off between convergence stability and model personalization, which often plagues traditional FL pipelines. The LPA component further ensures resilience to noisy or corrupted updates, particularly relevant for audio data subject to ambiguous labels, device noise, or inconsistent recording quality.
@@ 5 Convergence Analysis of FedMLAC
This section presents a rigorous convergence analysis of the FedMLAC framework, focusing on the global Plug-in model’s convergence under the challenges of data heterogeneity, model heterogeneity, and data poisoning. We formalize the theoretical assumptions, derive the convergence properties, and present a detailed proof to demonstrate that FedMLAC achieves convergence to a stationary point with a quantifiable rate.
@@ 5.2 Convergence Analysis
We analyze the convergence of the global Plug-in model $\theta^{g}$ to a stationary point of the global objective $\mathcal{L}(\theta^{g})$. The analysis proceeds in three steps: (1) bounding the loss decrease during local updates of $\omega^{k}$ and $\theta^{k}$, (2) analyzing the effect of LPA on global aggregation, and (3) deriving the overall convergence rate.
5.2.1 Local Update Analysis
For each client $k$, the local model $\omega^{k}$ is updated via SGD as below:
$\omega_{t+(i+1)}^{k}=\omega_{t+(i)}^{k}-\eta\nabla\mathcal{L}^{k}(\mathcal{B};\omega_{t+(i)}^{k},\theta_{t+(i)}^{k}),$ (21)
where $\eta$ is the learning rate, and $i=0,1,\ldots,E-1$ indexes local epochs. Since $\mathcal{L}^{k}$ is $L$-smooth (Assumption 1), we apply the descent lemma:
$\mathbb{E}[\mathcal{L}^{k}(\omega_{t+(i+1)}^{k})]$ $\leq\mathcal{L}^{k}(\omega_{t+(i)}^{k})-\eta\left\langle\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k}),\mathbb{E}\left[\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k};\mathcal{B})\right]\right\rangle$
$\quad+\frac{L\eta^{2}}{2}\mathbb{E}\left[\left\|\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k};\mathcal{B})\right\|^{2}\right].$ (22)
By Assumption 4, the stochastic gradient is unbiased: $\mathbb{E}[\nabla\mathcal{L}^{k}(\omega;\mathcal{B})]=\nabla\mathcal{L}^{k}(\omega)$. The gradient variance bound gives:
$\mathbb{E}[\|\nabla\mathcal{L}^{k}(\omega;\mathcal{B})\|^{2}]$ $=\mathbb{E}[\|\nabla\mathcal{L}^{k}(\omega;\mathcal{B})-\nabla\mathcal{L}^{k}(\omega)+\nabla\mathcal{L}^{k}(\omega)\|^{2}]$
$\leq\|\nabla\mathcal{L}^{k}(\omega)\|^{2}+\sigma_{g}^{2}$
$\leq G^{2}+\sigma_{g}^{2}.$ (23)
where the last inequality follows from Assumption 2. Thus:
$\mathbb{E}[\mathcal{L}^{k}(\omega_{t+(i+1)}^{k})]\leq\mathcal{L}^{k}(\omega_{t+(i)}^{k})-\eta\|\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}}{2}(G^{2}+\sigma_{g}^{2}).$ (24)
Summing over $E$ local epochs:
$\mathbb{E}[\mathcal{L}^{k}(\omega_{t+1}^{k})]\leq\mathcal{L}^{k}(\omega_{t}^{k})-\eta\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{k}(\omega_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}E}{2}(G^{2}+\sigma_{g}^{2}).$ (25)
Similarly, for the Plug-in model $\theta^{k}$, updated via:
$\theta_{t+(i+1)}^{k}=\theta_{t+(i)}^{k}-\eta\nabla\mathcal{L}^{g}(\mathcal{B};\omega_{t+(i)}^{k},\theta_{t+(i)}^{k}),$ (26)
we obtain:
$\mathbb{E}[\mathcal{L}^{g}(\theta_{t+(i+1)}^{k})]\leq\mathcal{L}^{g}(\theta_{t+(i)}^{k})-\eta\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}}{2}(G^{2}+\sigma_{g}^{2}).$ (27)
@@ 5.2.1 Local Update Analysis
$\mathbb{E}[\mathcal{L}^{g}(\theta_{t+1}^{k})]\leq\mathcal{L}^{g}(\theta_{t}^{k})-\eta\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}E}{2}(G^{2}+\sigma_{g}^{2}).$ (28)
The mutual learning mechanism, through the KL divergence terms in Eqs. (8) and  (7), aligns the local model $\omega^{k}$ with the global knowledge in $\theta^{k}$ and vice versa, reducing the impact of data heterogeneity by regularizing the local optimization direction.
5.2.2 Global Aggregation with LPA
In the FedMLAC-Aggregation step, the server aggregates the client-uploaded Plug-in models $\{\theta_{t+1}^{k}\}_{k\in S_{t}}$ using LPA. For each layer $l$, LPA computes the average parameters:
$\bar{\theta}_{t+1}^{l}=\frac{1}{|S_{t}|}\sum_{k\in S_{t}}\theta_{t+1}^{k,l},$ (29)
and the deviation for each client $k$:
$\delta^{k,l}=\|\theta_{t+1}^{k,l}-\bar{\theta}_{t+1}^{l}\|_{2}.$ (30)
The trusted subset $\mathcal{T}_{l}$ excludes the top $\lfloor v_{h}\cdot|S_{t}|\rfloor$ and bottom $\lfloor v_{l}\cdot|S_{t}|\rfloor$ clients based on $\delta^{k,l}$, and the global layer parameters are computed as:
where $n_{\mathcal{T}_{l}}=\sum_{k\in\mathcal{T}_{l}}n_{k}$. The global Plug-in model is $\theta_{t+1}^{g}=\{\theta_{t+1}^{g,l}\}_{l=1}^{L}$.
To analyze the effect of LPA, we consider the deviation of the aggregated model from the true global optimum. By Assumption 5, the parameters in $\mathcal{T}_{l}$ satisfy $\mathbb{E}[\|\theta_{t+1}^{k,l}-\bar{\theta}_{t+1}^{l}\|^{2}]\leq\delta^{2}$. The global loss after aggregation can be bounded using the smoothness of $\mathcal{L}$ (Assumption 1):
$\mathbb{E}[\mathcal{L}(\theta_{t+1}^{g})]\leq\mathcal{L}(\theta_{t}^{g})+\langle\nabla\mathcal{L}(\theta_{t}^{g}),\mathbb{E}[\theta_{t+1}^{g}-\theta_{t}^{g}]\rangle+\frac{L}{2}\mathbb{E}[\|\theta_{t+1}^{g}-\theta_{t}^{g}\|^{2}].$ (32)
The expected update direction is:
$\mathbb{E}[\theta_{t+1}^{g}-\theta_{t}^{g}]=\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}(\theta_{t+1}^{k,l}-\theta_{t}^{g,l}).$ (33)
Since $\theta_{t+1}^{k,l}$ results from $E$ local SGD steps, we approximate:
$\theta_{t+1}^{k,l}-\theta_{t}^{g,l}\approx-\eta\sum_{i=0}^{E-1}\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k};\mathcal{B}).$ (34)
Thus:
$\mathbb{E}[\theta_{t+1}^{g}-\theta_{t}^{g}]\approx-\eta\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k}).$ (35)
The second term, $\mathbb{E}[\|\theta_{t+1}^{g}-\theta_{t}^{g}\|^{2}]$, is influenced by the variance of local updates and the pruning effect of LPA:
$\mathbb{E}[\|\theta_{t+1}^{g}-\theta_{t}^{g}\|^{2}]\leq\eta^{2}E(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}),$ (36)
where $\sigma^{2}$ accounts for data heterogeneity (Assumption 3), and $\delta^{2}$ accounts for residual deviations after LPA (Assumption 5). Combining these, the global loss decrease is:
$\mathbb{E}[\mathcal{L}(\theta_{t+1}^{g})]\leq\mathcal{L}(\theta_{t}^{g})-\eta\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}+\frac{L\eta^{2}E}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (37)
5.2.3 Convergence Theorem
We now present the main convergence result for FedMLAC.
@@ 5.2.3 Convergence Theorem
Under Assumptions 16, with learning rate $\eta<\frac{1}{L}$, after $T$ communication rounds, the expected gradient norm of the global Plug-in model satisfies:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t}^{g})\|^{2}]\leq\frac{\mathcal{L}(\theta_{0}^{g})-\mathcal{L}^{*}}{\eta TE}+\frac{L\eta E}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (38)
Summing the global loss decrease (37) over $T$ rounds:
$\sum_{t=0}^{T-1}\mathbb{E}[\mathcal{L}(\theta_{t+1}^{g})-\mathcal{L}(\theta_{t}^{g})]$ $\leq-\eta\sum_{t=0}^{T-1}\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\left\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\right\|^{2}$
$\quad+\frac{L\eta^{2}ET}{2}\left(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}\right).$ (39)
By Assumption 6, the left-hand side is:
$\sum_{t=0}^{T-1}(\mathcal{L}(\theta_{t+1}^{g})-\mathcal{L}(\theta_{t}^{g}))=\mathcal{L}(\theta_{T}^{g})-\mathcal{L}(\theta_{0}^{g})\geq\mathcal{L}^{*}-\mathcal{L}(\theta_{0}^{g}).$ (40)
Rearranging:
$\eta\sum_{t=0}^{T-1}\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}\leq\mathcal{L}(\theta_{0}^{g})-\mathcal{L}^{*}+\frac{L\eta^{2}ET}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (41)
Assuming $\theta_{t+(i)}^{k}\approx\theta_{t}^{g}$ for small $\eta$, we approximate:
$\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta_{t+(i)}^{k})\|^{2}\approx E\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t}^{g})\|^{2}].$ (42)
Dividing by $\eta TE$, we obtain:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t}^{g})\|^{2}]\leq\frac{\mathcal{L}(\theta_{0}^{g})-\mathcal{L}^{*}}{\eta TE}+\frac{L\eta E}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (43)

### v2

@@ 3.2 FedMLAC
The second component, FedMLAC-Aggregation, operates on the server side. It aggregates the updated Plug-in AC models uploaded by the clients, resulting in an updated global Plug-in AC model. This updated global AC model serves as an auxiliary tool for the subsequent round of client AC model updates, as illustrated in Figure 1.
@@ 3.2.1 FedMLAC Update
3:  for each epoch $i$ from 111 to $E$ do
@@
5:     Compute loss $\mathcal{L}^{k}(\mathcal{B};\omega^{k}_{t},\theta_{t}^{k})$ using Equation (8)
@@
10:     Compute loss $\mathcal{L}^{g}(\mathcal{B};\omega^{k}_{t},\theta^{k}_{t})$ using Equation (9)
@@
Local AC Model Update
@@
wherein $KL(\cdot,\cdot)$ denotes the KL divergence loss. By combining Equations. (6) and (7), we reformulate the client AC model training objective function in the $(t+1)^{th}$ communication round as:
@@
Global Plug-in Model Update
@@
The loss of Equation (9) ensures that the Plug-in model effectively integrates client-specific AC data knowledge, enhancing its adaptability and robustness for global FL. The entire FedMLAC-Update process is detailed in Algorithm 1.
3.2.2 FedMLAC Aggregation
Once the global Plug-in model on each client completes its local knowledge update, it is uploaded to the server for aggregation to generate a new global Plug-in model. In traditional federated aggregation (e.g., Equation (3)), all client-uploaded Plug-in models $\{\theta_{t+1}^{k}\}_{k\in S_{t}}$ are directly combined using data-size weighted averaging. However, under non-IID data distributions or adversarial conditions (e.g., noisy inputs, label-flipping, or poisoned data), some local Plug-in models may introduce substantial deviations from the true global optimum. These deviations manifest as abnormally large or small parameters in specific layers, thereby biasing the global Plug-in model when such updates are indiscriminately aggregated.
@@ 3.2.2 FedMLAC Aggregation
This value reflects how much Plug-in model $k$’s $l$ parameters deviate from the average behavior of the cohort. A high deviation may indicate either statistical heterogeneity or an adversarial anomaly.
@@
4:   for each model layer $l$ from 111 to $L$ do
5:     Compute $\bar{\theta}_{t+1}^{l}$ using Equation (10)
@@
7:     Compute $\delta^{k,l}$ using Equation (11)
@@
12:  Sort and prune $\bar{\mathcal{T}}_{l}$ as $\mathcal{T}_{l}$ according to Equation (12)
@@
[…]
@@
3.3 Framework Insights
FedMLAC embodies a unified design that integrates personalization, scalability, and robustness in FedAC by decoupling global knowledge aggregation from personalized learning through a collaborative Plug-in mechanism. This separation supports model heterogeneity and allows clients to retain autonomy over model structures and local objectives, which is essential in diverse acoustic scenarios with varying tasks, devices, and environments. By framing mutual learning as bidirectional knowledge distillation, FedMLAC balances generalization and adaptation within each training round, mitigating the trade-off between convergence stability and personalization that often hinders traditional FL pipelines. The LPA component further enhances resilience to noisy or corrupted updates, which is critical for audio data affected by ambiguous labels, device distortions, or inconsistent recording quality.
@@ 4 Convergence Analysis of FedMLAC
This section presents a convergence analysis of the FedMLAC framework, focusing on the global Plug-in model’s convergence under the challenges of data heterogeneity, model heterogeneity, and data poisoning. We outline the theoretical assumptions, key analytical steps, and the main convergence result.
@@ 4.2 Convergence Analysis
We analyze the convergence of $\theta^{g}$ to a stationary point of $\mathcal{L}(\theta^{g})$ in three steps: (1) bounding local update progress, (2) analyzing LPA’s effect on aggregation, and (3) deriving the global convergence rate.
Clients update $\omega^{k}$ and $\theta^{k}$ via SGD on Equations (8) and (9). For $\omega^{k}$:
$\omega^{k}_{t+(i+1)}=\omega^{k}_{t+(i)}-\eta\nabla\mathcal{L}^{k}(\mathcal{B};\omega^{k}_{t+(i)},\theta^{k}_{t+(i)}),$ (22)
where $i=0,\dots,E-1$. By $L$-smoothness (Assumption 1), the expected loss decrease is:
$\mathbb{E}[\mathcal{L}^{k}(\omega^{k}_{t+(i+1)})]\leq\mathcal{L}^{k}(\omega^{k}_{t+(i)})-\eta\|\nabla\mathcal{L}^{k}(\omega^{k}_{t+(i)})\|^{2}+\frac{L\eta^{2}}{2}(G^{2}+\sigma_{g}^{2}).$ (23)
@@
$\mathbb{E}[\mathcal{L}^{k}(\omega^{k}_{t+1})]\leq\mathcal{L}^{k}(\omega^{k}_{t})-\eta\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{k}(\omega^{k}_{t+(i)})\|^{2}+\frac{L\eta^{2}E}{2}(G^{2}+\sigma_{g}^{2}).$ (24)
Similarly, for $\theta^{k}$:
$\mathbb{E}[\mathcal{L}^{g}(\theta^{k}_{t+1})]\leq\mathcal{L}^{g}(\theta^{k}_{t})-\eta\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta^{k}_{t+(i)})\|^{2}+\frac{L\eta^{2}E}{2}(G^{2}+\sigma_{g}^{2}).$ (25)
Equation (25) reflects the Plug-in model’s progress, driven by mutual learning with $\omega^{k}$. The KL divergence term in Equation (8) regularizes $\omega^{k}$ toward global knowledge, reducing the impact of data heterogeneity ($\sigma^{2}$).
In FedMLAC-Aggregation, LPA computes per-layer averages $\bar{\theta}^{l}_{t+1}$, prunes outliers based on deviations $\delta^{k,l}=\|\theta^{k,l}_{t+1}-\bar{\theta}^{l}_{t+1}\|_{2}$, and aggregates:
$\theta^{g,l}_{t+1}=\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\theta^{k,l}_{t+1}.$ (26)
Equation (26) forms the global model by weighting trusted client updates, with $\mathbb{E}[\|\theta^{k,l}_{t+1}-\bar{\theta}^{l}_{t+1}\|^{2}]\leq\delta^{2}$ (Assumption 5). The global loss decrease is:
$\mathbb{E}[\mathcal{L}(\theta^{g}_{t+1})]\leq\mathcal{L}(\theta^{g}_{t})-\eta\sum_{l=1}^{L}\sum_{k\in\mathcal{T}_{l}}\frac{n_{k}}{n_{\mathcal{T}_{l}}}\sum_{i=0}^{E-1}\|\nabla\mathcal{L}^{g}(\theta^{k}_{t+(i)})\|^{2}+\frac{L\eta^{2}E}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (27)
4.3 Convergence Theorem
@@
Consider the training dynamics of the global Plug-in model $\theta^{g}$ in FedMLAC under Assumptions 16. If the learning rate satisfies $\eta<\frac{1}{L}$, then after $T$ communication rounds, the average squared norm of the gradient of the global loss function $\mathcal{L}(\theta^{g})$ satisfies:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\|\nabla\mathcal{L}(\theta^{g}_{t})\|^{2}\right]\leq\frac{\mathcal{L}(\theta^{g}_{0})-\mathcal{L}^{*}}{\eta TE}+\frac{L\eta E}{2}(G^{2}+\sigma_{g}^{2}+\sigma^{2}+\delta^{2}).$ (28)
Equation (28) establishes that the global Plug-in model converges to a stationary point at a rate of $\mathcal{O}(1/\sqrt{T})$ (with $\eta\propto 1/\sqrt{T}$). The bound captures error contributions from gradient magnitude ($G^{2}$), local stochastic variance ($\sigma_{g}^{2}$), data heterogeneity ($\sigma^{2}$), and deviations induced by LPA pruning ($\delta^{2}$).